OpenStudy (anonymous):

help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! (−3 + 2i) • (2 + i)

OpenStudy (anonymous):

The order of operations holds for complex numbers as it does for reals. Thus (-3 + 2i) * (2 + i) = ( -2 + 4i - 3i -6 ) = i - 8